OpenStudy (anonymous):

Solve the following system. y = x + 3 2x + y = 9 (2, 5) (5, 2) (−2, 5) (2, −5)

OpenStudy (anonymous):

take the points and plug in for x and y 5=2+3 2(2)+5=9

OpenStudy (anonymous):

i think its B but im not sure

OpenStudy (anonymous):

the must be =

OpenStudy (anonymous):

its not B

OpenStudy (anonymous):

or you can do it in different way

OpenStudy (anonymous):

2x+x+3=9 3x=6 x=2 and plug in x and find the y

OpenStudy (anonymous):

@mikaa_toxica13

OpenStudy (anonymous):

not quite sure what to do with it now

OpenStudy (anonymous):

so x is 2 right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so take one of the equation and for x put 2 so y=2+3 and you will find y

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

can you show the work please

OpenStudy (anonymous):

im not too sure its right. my teacher explained it too me and she kept repeating -5

OpenStudy (anonymous):

but the x is 2 so to fin do solve this equation y=2+3 y=

OpenStudy (anonymous):

couldn't be -5

OpenStudy (anonymous):

it would be 5 ! -_- im retarded lol

OpenStudy (anonymous):

yes, so x=2 and y=5 no you are not :)

OpenStudy (anonymous):

so (2, 5)

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

thank u !!!

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

can u help with another ? its simpler

OpenStudy (anonymous):

@begzat

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Part 1: For the following system of equations, write your own real world scenario that describes what is happening. Use complete sentences and correct grammar in your scenario. Part 2: Solve the system and explain what the results mean according to your scenario. 2x + y = 9 3x + 4y = 26

OpenStudy (anonymous):

or u can go on my new question lol